If FG = x + 3, GH = 4x, and FH = 8, what is FG?
Found 2 solutions by ikleyn, mananth:
Answer by ikleyn(52797)

.

As it is worded, printed and presented in the post, the problem is INCOMPLETE.

To be complete, it should say that the point G lies on the straight line between
the points F and H. I will assume it, by default.

Then FG + GH = FH, or

    (x+3) + (4x) = 8.


It implies

    5x = 8 - 3

    5x = 5

     x = 5/5 = 1

    FG = x+3 = 1 + 3 = 4.


ANSWER.  Under the accepted assumption, FG = 4.

Solved.



Answer by mananth(16946)

If FG = x + 3, GH = 4x, and FH = 8, what is FG?

4x+x+3=8
5x+3=8
subtract 3 from both sides
5x = 5
x=1
FG = x+3= 1+3=4